An unknown resistor is connected between the terminals of a 6.5 V battery. Energy is dissipated in the resistor at the rate of 0.480 W. The same resistor is then connected between the terminals of a 3.58 V battery. At what rate is energy now dissipated?
Answer #1

Resistance

R=V2/P =6.52/0.48

R=88.02 ohms

Energy is dissipated at a rate when it is connected to V=3.58 Volts

P=V2/R=3.582/88.02

P=0.1456 Watts
Answer #3

P = V^2/R

R =V^2/P

R = 6.6*6.5/.480 = 88.02 ohm

P = V^2/R

P = 0.146 W
Answer #4

Energy is dissipated in the resistor at the rate=power loss across resistor=v*v/r